OpenStudy (anonymous):

can someone please explain what a quadratic formula is and how to solve equations using it

OpenStudy (anonymous):

The quadratic formula is a general solution for all equations of the form \(ax^2+bx+c=0\). This means that it is an expression for \(x\) in terms of \(a,b,\text{and } c\)...essentially it tells you how to solve any quadratic equation.

OpenStudy (anonymous):

Its definition is\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] So suppose you wanted to solve \(2x^2-5x-6\) using the quadratic formula. Then you would simply plug in \(a=2\), \(b=-5\), and \(c=-6\) and out would come your answer.

OpenStudy (anonymous):

ok so would that work in a equation like 9n^2 =4+7n

OpenStudy (anonymous):

Yep. You just need to get it into the form \(an^2+bn+c=0\)...so you would have \(9n^2-7n-4=0\), and then you can plug it in.

OpenStudy (anonymous):

i still dont quite get it what are you plugging in

OpenStudy (anonymous):

You are letting \(a=9\), \(b=-7\), and \(c=-4\) (the coefficients), and then you are taking \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] and replacing every \(a\) with 9, every \(b\) with -7, every \(c\) with -4, and then evaluation, which will give you \(x\), which will be your answers (note the \(\pm\)).

OpenStudy (anonymous):

wait what happens to the n's

OpenStudy (anonymous):

You're solving for n.

OpenStudy (anonymous):

ok thanks